# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

### v3

@@ Abstract.
The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra. Masaki Kameko [3] constructed, in a distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a decomposition in which the top layer is spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates. While Kameko obtained a restriction on the possible $Q_{0}$–image in the top layer, we determine the corresponding hit image in $V$ exactly. We make three main contributions as follows:
First, we give an exact structural description of the hit subspace on the top-layer summand $V=\langle\sigma(z_{k})\rangle$ by constructing a parity functional $\varepsilon:V\to\mathbb{F}_{2}$ and proving that $V\cap(\text{hit subspace})=\ker(\varepsilon)$; equivalently, parity yields a complete criterion for hitness in the $M_{1}$–summand and $V/(\text{hits})\cong\mathbb{F}_{2}$. In particular, this computes the quotient $V/(\text{hits})$. We next explain how this parity classification interacts with the numerical condition $\beta(d)>n$. In particular, the counterexamples arise as consequences of the structural theorem.
Second, we combine this classification with the numerical condition $\beta(d)>n$ to obtain a new infinite counterexample family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$), distinct from Kameko’s k=n3k=n-3 family.
Third, we prove a base-change invariance statement: the parity classification on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
@@ 1. Introduction
Kameko’s proof is quite delicate. He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer of the motivic hit quotient. He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
The present paper strengthens Kameko’s analysis of the top layer from an inclusion statement to an exact structural description. In Kameko’s decomposition, the summand
$V=\big\langle\sigma(z_{k})\ \big|\ \sigma\in\mathrm{Mono}(k)\big\rangle$
captures the entire “top layer” contribution in degree $d=k+2d_{1}$. Kameko showed that only certain linear combinations can arise from the Bockstein $Q_{0}$ in this layer, namely combinations generated by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$, yielding a containment for the possible $Q_{0}$–image in $V$. What remained open is whether this containment is sharp, i.e. whether it actually exhausts the hit image inside $V$, and hence what the precise obstruction to being hit is. We answer this question by giving a complete parity-based description of the hit image in $V$.
Accordingly, our main contribution in this work is structural: we upgrade Kameko’s top-layer restriction from a containment to an equality by giving an exact parity-based description of the hit image inside $V$. In particular, the counterexample families arise as consequences of this classification, once combined with arithmetic input.
@@
    Parity classification (exact hit image). We construct an explicit linear functional $\varepsilon:V\to\mathbb{F}_{2}$ (the parity) and prove an if-and-only-if criterion: a class in $V$ is hit precisely when it has even parity. Equivalently, the hit image in $V$ is the codimension-one hyperplane $\ker(\varepsilon)$, so that $V/(\text{hits})\cong\mathbb{F}_{2}$.
@@
    Arithmetic application (new counterexample family). As a consequence, once $\beta(d)>n$ holds in these degrees, any odd-parity class in $V$ yields a counterexample to the motivic Peterson-type conjecture. Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop. 2.1], we establish a new infinite family with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) for which $\beta(d)>n$.
@@
    Base change. We show that the parity classification on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
Our main structural result is the following, which establishes that parity is the precise obstruction to being hit in the summand $V$.
Theorem 1.4 (Parity classification in the top layer).
For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span in $QN_{n}^{d,*}$ of the monotone translates $\sigma(z_{k})$ with $\sigma\in\mathrm{Mono}(k)$. Let $\theta:N_{n}^{d,*}\to V$ be the projection onto the $M_{1}$–component (defined in Section 3).
There exists a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
@@
    Exact hit image in $V$:
    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
@@
    If and only if criterion: For $v\in V$, one has $v$ is hit in $QN_{n}^{d,*}$ if and only if $\varepsilon(v)=0$. In particular $V/(\text{hits})\cong\mathbb{F}_{2}$, generated by any $v\in V$ with $\varepsilon(v)=1$.
Remark 1.5 (Sharpness).
Theorem 1.4 upgrades Kameko’s restriction [3] on the possible $Q_{0}$–image in the top layer to an equality statement: the hit image inside $V$ is a single explicit codimension-one hyperplane. Thus the obstruction in this summand is minimal (one-dimensional) and is completely detected by parity. Equivalently, the theorem computes the one-dimensional quotient $V/(\text{hits})$, and provides an effective parity criterion: a class in $V$ is hit if and only if it has even parity.
On the arithmetic side, we prove:
Theorem 1.6 (New numerical family with $\beta(d)>n$).
Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let
$d=(n-1)\bigl(2^{k+1}-2\bigr)+k$
be the degree of $z_{k}$. Then $\beta(d)>n$.
Combining Theorems 1.3, 1.4 and 1.6 with the base-change discussion in Section 5, we obtain the following.
@@
    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}^{d,*}(K)$ with $\sigma\in\mathrm{Mono}(k)$ represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus, combining the parity description of the top-layer hit image with the numerical condition $\beta(d)>n$, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
The paper is organized as follows. In Section 2 we review the classical hit problem, the motivic Steenrod algebra, and Kameko’s description of the motivic cohomology of $BV_{n}$ and the motivic hit problem. Section 3 introduces the parity functional on the top layer of the motivic hit quotient and proves Theorem 1.4. Section 4 is devoted to the arithmetic of the degrees $d=(n-1)(2^{k+1}-2)+k$, including Kameko’s original estimates and the new family in Theorem 1.6. Finally, Section 5 discusses the extension to arbitrary algebraically closed base fields of characteristic 0 and proves Theorem 1.7.
@@ 2.3. The motivic hit problem and the monomials $z_{k}$
The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module, and there is an isomorphism
@@
The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer of the hit quotient in terms of a pair of sets of monomials $M_{0}$ and $M_{1}$ which we review next.
@@ 2.4. Weight vectors and the subspaces $F_{n}$, $G_{n}$, and $H_{n}$
in the lexicographic order. He shows that the quotient $N_{n}/H_{n}$ decomposes as a direct sum over bigraded components $\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b}$ with $a+2b=d$ and an explicit description of the contribution of $Q_{0}$ [3, Prop. 4.1].
We will only need the degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis. Moreover, he proves that
@@
3. A parity functional on Kameko’s top layer
In this section we work in the situation of the previous section with degree $d=k+2d_{1}$, $d_{1}=(n-1)(2^{k}-1)$, and monomial $z_{k}$ as in (2.2). We define a linear functional on the $\mathbb{F}_{2}$–span of the monomials in $M_{1}$ and show that its kernel exactly describes the intersection with the image of the motivic Steenrod algebra.
3.1. The $M_{1}$–summand and its parity functional
Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Following Kameko, we consider the summand
$\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}\subset(N_{n}/H_{n})_{d,*},$
and the sets $M_{0},M_{1}\subset\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ whose images form a basis of this summand. By [3, Eq. (4.3)], we have
(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\},$
and the images of $M_{0}\cup M_{1}$ in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis.
Let $\pi:Y_{n}\to Y_{n}/G_{n}$ be the quotient map, and extend it to
$\pi:\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}\to\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
@@ 3.1. The $M_{1}$–summand and its parity functional
$V:=\Bigl\langle\pi\bigl(\sigma(z_{k})\bigr)\,\Big|\,\sigma\in\mathrm{Mono}(k)\Bigr\rangle\subset\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Since the $\pi(\sigma(z_{k}))$ are part of a basis, they are linearly independent and form a basis of $V$.
Definition 3.1 (Parity functional on the $M_{1}$–summand).
@@
Thus $\varepsilon(v)$ simply records whether the number of basis vectors $\pi(m_{i})$ that appear with coefficient $1$ in $v$ is even or odd.

### v4

@@ Abstract.
The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra. Masaki Kameko [3] constructed, in the distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a top layer spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates. In this paper, we work on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the hit subspace. We therefore construct a local projection
$\vartheta:N_{n}^{d,*}\longrightarrow V,$
from the raw degree–$d$ component to Kameko’s $M_{1}$–summand $V=\langle\{\sigma(z_{k})|\ \sigma\in\mathrm{Mono}(k)\}\rangle$, together with a parity functional $\varepsilon:V\to\mathbb{F}_{2}$, and prove that
$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
Thus parity gives an exact description of the local top-layer image of hit elements. As a consequence, any element whose local top-layer component has odd parity is non-hit. In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the motivic hit quotient.
Subsequently, a direct binary calculation shows that for
$n=2^{r}+1,\qquad k=n-4,\qquad r\geq 5,$
one has $\beta(d)>n$ for
$d=(n-1)(2^{k+1}-2)+k.$
Combined with Kameko’s non-hit theorem, this yields a new infinite family of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s k=n3k=n-3 family. Our local parity criterion sharpens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
Finally, we prove a base-change invariance statement: the local parity criterion on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
@@ 1. Introduction
Kameko’s proof is quite delicate. He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer in the relevant filtration. He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
The present paper strengthens Kameko’s top-layer analysis by working on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the full hit subspace. The key point is that, in degree $d=k+2d_{1}$, the part detected by Kameko’s set
$V=\bigl\langle\pi(\sigma(z_{k}))\ \big|\ \sigma\in\mathrm{Mono}(k)\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
should be viewed as a local top layer, rather than as a subspace of the final quotient by hits. Here and throughout, $\mathrm{Mono}(k)$ denotes the set of monotone injections
$\sigma:\{1,\dots,k\}\hookrightarrow\{1,\dots,n\}.$
Equivalently, $\mathrm{Mono}(k)$ may be identified with the set of kk–subsets of $\{1,\dots,n\}$ via $\sigma\mapsto\mathrm{Im}(\sigma)$. Once this distinction is made, Kameko’s containment [3] for the $Q_{0}$–image becomes an exact parity statement for the local image of hit elements.
Our main contribution is therefore structural: we construct a local projection
$\vartheta:N_{n}^{d,*}\longrightarrow V$
and an explicit parity functional $\varepsilon:V\to\mathbb{F}_{2}$ such that
$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
This gives a complete description of which vectors occur as the $M_{1}$–part of hit elements, and it yields an effective odd-parity obstruction.
@@
    Local parity criterion. We prove that, in degree $d=k+2d_{1}$, the local top-layer image of hit elements is exactly the codimension-one hyperplane $\ker(\varepsilon)\subset V$. Equivalently, an element of $N_{n}^{d,*}$ whose local top-layer component has odd parity cannot be hit. In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit.
@@
    Arithmetic consequence and refinement. A direct binary calculation shows that if $n=2^{r}+1$ and k=n4k=n-4 with $r\geq 5$, then $\beta(d)>n$ for
    $d=(n-1)(2^{k+1}-2)+k.$
    Combined with Kameko’s non-hit theorem [3], this gives a new infinite family of motivic Peterson counterexamples. The local parity criterion refines this by proving that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
@@
    Base change. We show that the local parity criterion on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
Our main structural result is the following.
Theorem 1.4 (Local parity criterion in the top layer).
For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span of the monotone translates $\sigma(z_{k})$ in Kameko’s $M_{1}$–summand. There exist a linear map
$\vartheta:N_{n}^{d,*}\longrightarrow V$
and a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
@@
    Exact local top-layer hit image:
    $\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
@@
    Odd-parity obstruction: if $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A_{+}^{\sharp}(N_{n})$.
  3. (iii)
    In particular, any odd-parity linear combination of the monomials $\sigma(z_{k})$ is non-hit and therefore represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and in $QM_{n}^{d,*}(\mathbb{C})$.
Remark 1.5 (Why the local formulation matters).
The map $\vartheta$ is defined on the raw degree–$d$ piece $N_{n}^{d,*}$, before quotienting by $A_{+}^{\sharp}(N_{n})$. This is essential: if one first passes to a quotient that already modds out hit elements, then every hit class becomes zero and the top-layer information is lost. The theorem above isolates the correct statement: parity detects exactly the local top-layer parts of hit elements.
On the arithmetic side, we show by a direct binary calculation that for $r\geq 5$, the choice $n=2^{r}+1$ and k=n4k=n-4 satisfies $\beta(d)>n$ for
$d=(n-1)(2^{k+1}-2)+k.$
Together with Kameko’s non-hit theorem, this yields a new infinite family of motivic Peterson counterexamples, distinct from Kameko’s original case k=n3k=n-3. The local parity criterion strengthens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the corresponding motivic hit quotient.
Combining Theorem 1.3, the local parity criterion of Theorem 1.4, the direct binary calculation for the case $n=2^{r}+1$, k=n4k=n-4, and the base-change discussion in Section 4, we obtain the following.
@@
    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}(K)_{d,*}$, with $\sigma\in\mathrm{Mono}(k)$, represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus, combining the local parity description of the top layer with the direct verification of the inequality $\beta(d)>n$ in the case $n=2^{r}+1$, k=n4k=n-4, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
The paper is organized as follows. In Section 2 we review the classical hit problem, the motivic Steenrod algebra, and Kameko’s description of the motivic cohomology of $BV_{n}$ and the motivic hit problem. Section 3 introduces the local top-layer projection, defines the parity functional on Kameko’s $M_{1}$–summand, and proves Theorem 1.4. Finally, Section 4 discusses the extension to arbitrary algebraically closed base fields of characteristic 0, proves the base-change invariance of the local parity criterion, and combines it with the relevant arithmetic input to establish Theorem 1.6.
@@ 2.3. The motivic hit problem and the monomials $z_{k}$
Lemma 2.3 (Passage modulo $\tau$ preserves hit elements).
Let
$q_{\tau}:M_{n}(K)\longrightarrow N_{n}(K)=M_{n}(K)/(\tau)$
be the quotient map. If
$u\in A^{>0,*}\cdot M_{n}(K)\cap M_{n}(K)_{d,*},$
then
$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K))\cap N_{n}(K)_{d,*}.$
Consequently, $q_{\tau}$ induces a surjective linear map
$QM_{n}^{d,*}(K)\longrightarrow QN_{n}^{d,*}(K).$
Proof.
Write
$u=\sum_{i}a_{i}\cdot v_{i}\qquad(a_{i}\in A^{>0,*},\ v_{i}\in M_{n}(K)).$
Passing to the quotient modulo $\tau$, we obtain
$q_{\tau}(u)=\sum_{i}\bar{a}_{i}\cdot q_{\tau}(v_{i}),$
where $\bar{a}_{i}$ denotes the image of $a_{i}$ in
$A^{\sharp}=A^{*,*}/(\tau).$
If $\bar{a}_{i}=0$, the corresponding term vanishes. If $\bar{a}_{i}\neq 0$, then $\bar{a}_{i}$ still has positive total degree, hence $\bar{a}_{i}\in A^{\sharp}_{+}$. Therefore
$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K)).$
This proves the first assertion. The induced surjection on hit quotients follows immediately from the definition. ∎
The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module. By Lemma 2.3, the quotient map modulo $\tau$ induces a surjective map on hit quotients; in fact, one has an isomorphism
@@
The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer in terms of two sets of monomials $M_{0}$ and $M_{1}$ which we review next.
@@ 2.4. Weight vectors and the subspaces $F_{n}$, $G_{n}$, and $H_{n}$
in the lexicographic order. He shows that the quotient $(N_{n}/H_{n})_{d,*}$ decomposes as a direct sum of pieces
$\Lambda_{n}^{a}\otimes(Y_{n}/G_{n})^{2b}/\pi\bigl(Q_{0}(\Lambda_{n}^{a+1}\otimes Y_{n}^{2(b-1)})\bigr)\qquad(a+2b=d),$
see [3, Sec. 4]. We will only need the degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis. Moreover, he proves that
@@
3. A parity functional on Kameko’s local top layer
In this section we work in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Our goal is to define a projection onto the $M_{1}$–summand before quotienting by hit elements and then to identify the image of hit elements under that projection.
3.1. The local top-layer map and the parity functional
Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Since
$N_{n}^{d,*}=\bigoplus_{a+2b=d}\Lambda_{n}^{a}\otimes Y_{n}^{2b},$
there is a canonical projection
$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$
onto the summand with exactly $k$ exterior generators.
Let
$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
be the quotient map induced by $Y_{n}\to Y_{n}/G_{n}$. By Kameko’s description recalled above, there are subsets $M_{0},M_{1}\subset\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ such that $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ and
(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\}.$
@@ 3.1. The local top-layer map and the parity functional
$U_{0}:=\langle\pi(m)\mid m\in M_{0}\rangle,\qquad V:=\langle\pi(m)\mid m\in M_{1}\rangle.$
Then
$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
Let
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
be the projection with kernel $U_{0}$.
Definition 3.1 (Local top-layer projection).
Define the linear map
$\vartheta:=p_{M_{1}}\circ\pi\circ\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow V.$
We call $\vartheta(u)$ the local $M_{1}$–component (or local top-layer component) of $u\in N_{n}^{d,*}$.
The point is that $\vartheta$ is defined directly on $N_{n}^{d,*}$. It records the $M_{1}$–part of an element after discarding only the lower-weight terms encoded by $G_{n}$ and the complementary basis block $M_{0}$; it does not mod out by hit elements.
Definition 3.2 (Parity functional on the local top layer).
@@
Thus $\varepsilon(v)$ records whether the number of basis vectors $\pi(m_{i})$ appearing in $v$ is even or odd.